OpenStudy (anonymous):

Find the zeros of the function. f(x) = 8x3 + 40x2 + 48x PLEASE HELP!! MEDAL!!!!!!!

OpenStudy (anonymous):

You can do it in many ways...If you know calculus, then just differentiate the function...the thing that will be left....factorize it and find out the values of x for which the function will give zero...

OpenStudy (solomonzelman):

set f(x)=0, factor out of 8x and then solve the quadratic by factoring

OpenStudy (anonymous):

0 is one right?

OpenStudy (anonymous):

so are the zeros 0,2, and 3?

OpenStudy (anonymous):

@SolomonZelman am i right?

OpenStudy (solomonzelman):

\(\large\color{black}{ \bf f(x) = 8x^3 + 40x^2 + 48x}\) \(\large\color{black}{ \bf 8x^3 + 40x^2 + 48x=0}\) \(\large\color{black}{ \bf 8x(x^2 + 5x + 6x)=0~~~~~~~~~~8x=0~~->~~x=0,}\) \(\large\color{black}{ \bf x^2 + 5x + 6x=0}\) \(\large\color{black}{ \bf (x+2)(x+3)=0~~~~~~~~~~x=~-2,~~-3~}\)

OpenStudy (anonymous):

8x(x^2+5x+6)

Parth (parthkohli):

@SolomonZelman I liked that answer. It was very \(\mathbb{bold}\)ly given. :D

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

so the zeros are 0, -2, and -3? @SolomonZelman

OpenStudy (anonymous):

@SolomonZelman are those the zeros?

OpenStudy (solomonzelman):

yes